# Mirror and Preconditioned Gradient Descent in Wasserstein Space

### v1

@@ 5 Applications and Experiments
In this section, we first discuss how to verify the relative convexity and smoothness between functionals in practice. Then, we provide some examples of mirror descent and preconditioned gradient descent on different objectives. We refer to Appendix G for more details on the experiments.
@@
To assess relative convexity or smoothness as stated in Definition 3, we need to compare the Bregman divergences along the right curves. When both functionals are of the same type, for example potential (respectively interaction) energies, this property is lifted from the convexity and smoothness on $\mathbb{R}^{d}$ of the underlying potential functions (respectively interaction kernels) to $\mathcal{P}_{2}(\mathbb{R}^{d})$, see Section E.2 for more details. When both are potential energies, the schemes (4) and (10) are equivalent to parallel MD and preconditioned GD since there are no interactions between the particles, and the conditions of convergences coincide with the ones obtained for MD and preconditioned GD on $\mathbb{R}^{d}$. In other cases, this provide schemes that are novel to the best of our knowledge. For functionals which are not of the same type, it is less straightforward. Using equivalent notions of convexity (Proposition 11), we may instead compare their Hessians along the right curves, see Section E.2 for an example between an interaction and a potential energy. For a functional obtained as a sum $\mathcal{F}=\mathcal{G}+\mathcal{H}$ with $\tilde{\mathcal{G}}_{\mu}$ and ~$\tilde{\mathcal{H}}_{\mu}$ convex, since $\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}}=\mathrm{d}_{\tilde{\mathcal{G}}_{\mu}}+\mathrm{d}_{\tilde{\mathcal{H}}_{\mu}}$, $\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}}\geq\max\{\mathrm{d}_{\tilde{\mathcal{G}}_{\mu}},\mathrm{d}_{\tilde{\mathcal{H}}_{\mu}}\}$, and thus $\mathcal{F}$ is 1-convex relative to $\mathcal{G}$ and $\mathcal{H}$. This includes e.g. the KL divergence which is convex relative to the potential and the negative entropy.
@@
MD on interaction energies.
We first focus on minimizing interaction energies $\mathcal{W}$ with kernel $W(z)=\frac{1}{4}\|z\|_{\Sigma^{-1}}^{4}-\frac{1}{2}\|z\|_{\Sigma^{-1}}^{2}$ with $\Sigma\in S_{d}^{++}(\mathbb{R})$, whose minimizer is an ellipsoid [26]. Since its Hessian norm can be bounded by a polynomial of degree 2, following [81, Section 2], $W$ is smooth relative to $K_{4}(z)=\frac{1}{4}\|z\|_{2}^{4}+\frac{1}{2}\|z\|_{2}^{2}$ and $\mathcal{W}$ is smooth relative to $\phi_{\mu}(\mathrm{T})=\iint K_{4}\big{(}\mathrm{T}(x)-\mathrm{T}(y)\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(y)$. Supposing additionally that the distributions are compactly supported, we can show that $\mathcal{W}$ is smooth relative to the interaction energy with $K_{2}(z)=\frac{1}{2}\|z\|_{2}^{2}$. For ill-conditioned $\Sigma$, the convergence can be slow. Thus, we also propose to use $K_{2}^{\Sigma}(z)=\frac{1}{2}\|z\|_{\Sigma^{-1}}^{2}$ and $K_{4}^{\Sigma}(z)=\frac{1}{4}\|z\|_{\Sigma^{-1}}^{4}+\frac{1}{2}\|z\|_{\Sigma^{-1}}^{2}$. We illustrate these schemes on Figure 2 and observe the convergence we expect for the schemes taking into account $\Sigma$. In practice, since $\nabla\phi_{\mu}(\mathrm{T})=(\nabla K\star\mathrm{T}_{\#}\mu)\circ\mathrm{T}$, the scheme needs to be approximated using Newton’s algorithm which can be computationally heavy. Using $\phi_{\mu}^{V}(\mathrm{T})=\int V\circ\mathrm{T}\ \mathrm{d}\mu$ with $V=K_{2}^{\Sigma}$, we obtain a more computationally friendly scheme with the same convergence, see Section G.2, but for which the smoothness is trickier to show.
MD on KL.
We now focus on minimizing $\mathcal{F}(\mu)=\int V\mathrm{d}\mu+\mathcal{H}(\mu)$ for $V(x)=\frac{1}{2}x^{T}\Sigma^{-1}x$ with $\Sigma$ possibly ill-conditioned, whose minimizer is the Gaussian $\nu=\mathcal{N}(0,\Sigma)$, and for which $\mathrm{W}_{2}$-gradient descent is slow to converge. We study the MD scheme in (4) with negative entropy $\mathcal{H}$ as the Bregman potential (NEM), and compare it on Figure 2 with the Forward-Backward (FB) scheme studied in [40] and the ideally preconditioned Forward-Backward scheme (PFB) with Bregman potential $\phi_{\mu}^{V}$ (see (115) in Appendix F). For computational purpose, we restrain the minimization in (4) over affine maps, which can be seen as taking the gradient over the submanifold of Gaussians [68, 40]. Starting from $\mathcal{N}(0,\Sigma_{0})$, the distributions stay Gaussian over the flow, and their closed-form is reported in (63) (Section D.3). We note that this might not be the case for the scheme (4), and thus that this scheme does not enter into the framework developed in the previous sections. Nonetheless, it demonstrates the benefits of using different Bregman potentials. We generate 20 Gaussian targets $\nu$ on $\mathbb{R}^{10}$ with $\Sigma=UDU^{T}$, $D$ diagonal and scaled in log space between 1 and 100, and $U$ a uniformly sampled orthogonal matrices, and we report the averaged KL over time. Surprisingly, NEM, which does not require an ideal (and not available in general) preconditioner, is almost as fast to converge as the ideal PFB, and much faster than the FB scheme.
@@
Preconditioned GD for single-cells.
Predicting the response of cells to a perturbation is a central question in biology. In this context, as the measuring process is destructive, feature descriptions of control and treated cells must be dealt with as (unpaired) source $\mu$ and target distributions $\nu$. Following [104], OT theory to recover a mapping $T$ between these two populations has been used in [20, 22, 21, 36, 45, 112, 64]. Inspired by the recent success of iterative refinement in generative modeling, through diffusion [57, 107] or flow-based models [78, 74], our scheme (1) follows the idea of transporting $\mu$ to $\nu$ via successive and dynamic displacements instead of, directly, with a static map $\bar{T}$. We model the transition from unperturbed to perturbed states through the (preconditioned) gradient flow of a functional $\mathcal{F}(\mu)=D(\mu,\nu)$ initialized at $\mu_{0}=\mu$, where $D$ is a distributional metric, and predict the perturbed population via $\hat{\mu}=\min_{\mu}\mathcal{F}(\mu)$. We focus on the datasets used in [20], consisting of cell lines analyzed using (i) 4i [54], and (ii) scRNA sequencing [110]. For each profiling technology, the response to respectively (i) 34 and (ii) 9 treatments are provided. As in [20], training is performed in data space for the 4i data and in a latent space learned by the scGen autoencoder [80] for the scRNA data. We use three metrics: the Sliced-Wasserstein distance $\mathrm{SW}_{2}^{2}$ [16], the Sinkhorn divergence $\mathrm{S}_{\varepsilon,2}^{2}$ [47] and the energy distance $\mathrm{ED}$ [98, 56, 55], and we compare the performances when minimizing this functional via preconditioned GD vs. (vanilla) GD. We measure the convergence speed when using a fixed relative tolerance $\textrm{tol}=10^{-3}$, as well as the attained optimal value $\mathcal{F}(\hat{\mu})$. Note that we follow [20] and additionally consider 40% of unseen (test) target cells for evaluation, i.e., for computing $\mathcal{F}(\hat{\mu})=D(\hat{\mu},\nu)$. As preconditioner, we use the one induced by $h^{*}(\*x)=(\|x\|_{2}^{a}+1)^{1/a}-1$ with $a>0$, which is well suited to minimize functionals which grow in $\|x-x^{*}\|^{a/(a-1)}$ near their minimum [111]. We set the step size $\tau=1$ for all the experiments. Then, we tune $a$ very simply: for a given metric $D$ and a profiling technology, we pick a random treatment and select $a\in\{1.25,1.5,1.75\}$ by grid search, and we generalize the selected $a$ for all the other treatments. Results are described in Figure 3: Preconditioned GD significantly outperforms GD over the 43 datasets, in terms of convergence speed and optimal value $\mathcal{F}(\hat{\mu})$. For instance, for $D=\mathrm{S}_{2,\varepsilon}^{2}$, we converge in 10 times less iterations while providing, on average, a better estimate of the treated population. We also compare our iterative (non parametric) approach with the use of a static (non parametric) map in Section G.4.
@@ 6 Conclusion
Clément Bonet acknowledges the support of the center Hi! PARIS. Adam David gratefully acknowledges funding by the BMBF 01|S20053B project SALE. Pierre-Cyril Aubin-Frankowski was funded by the FWF project P 36344-N. Anna Korba acknowledges the support of ANR-22-CE23-0030.
@@ Appendix A Related works
The space of probability distributions can be endowed with different metrics. When endowed with the Fisher-Rao metric instead of the Wasserstein distance, the geometry becomes very different. Notably, the shortest path between the two distributions is now a mixture between them. In this situation, the gradient is the first variation. Aubin-Frankowski et al. [8] studied the mirror descent in this space and notably showed connections with Sinkhorn algorithm when the mirror map and the optimized function are KL divergences. Karimi et al. [62] extended the mirror descent algorithm for more general time steps, and notably recovered the “Wasserstein Mirror Flow” proposed by Deb et al. [38] as a special case.
@@ B.1 Differential calculus on $L^{2}(\mu)$
The Fréchet derivative of $\phi$ denoted $\delta\phi$ is defined implicitly by
@@
If $\phi$ is Fréchet differentiable, then it is also Gâteaux differentiable, and both derivatives agree, i.e. for all $\mathrm{T},h\in L^{2}(\mu)$, $\delta\phi(\mathrm{T},h)=\phi^{\prime}(\mathrm{T})(h)$ [92, Proposition 1.26].
@@ B.2 Convexity on $L^{2}(\mu)$
which is equivalent by [92, Proposition 3.10] with
@@
We now present equivalent definitions of the relative smoothness and relative convexity, which is the equivalent of [81, Proposition 1.1].
@@
    $\psi$ $\beta$-smooth relative to $\phi$
@@
    $\psi$ $\alpha$-convex relative to $\phi$
@@
For the rest of the equivalences, we apply [92, Proposition 3.10]. Indeed, $\beta\phi-\psi$ convex is equivalent with
$\forall\mathrm{T},\mathrm{S}\in L^{2}(\mu),\ \langle\nabla(\beta\phi-\psi)(\mathrm{T})-\nabla(\beta\phi-\psi)(\mathrm{S}),\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)}\geq 0\\ \iff\forall\mathrm{T},\mathrm{S}\in L^{2}(\mu),\ \beta\langle\phi(\mathrm{T})-\nabla\phi(\mathrm{S}),\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)}\geq\langle\nabla\psi(\mathrm{T})-\nabla\psi(\mathrm{S}),\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)},$ (21)
which gives the equivalence between (a2) and (a4). And if $\psi$ and $\phi$ are twice differentiables, it is also equivalent with
$\forall\mathrm{T},\mathrm{S}\in L^{2}(\mu),\ \langle\nabla^{2}(\beta\phi-\psi)(\mathrm{T})\mathrm{S},\mathrm{S}\rangle_{L^{2}(\mu)}\geq 0\\ \iff\forall\mathrm{T},\mathrm{S}\in L^{2}(\mu),\ \beta\langle\nabla^{2}\phi(\mathrm{T})\mathrm{S},\mathrm{S}\rangle_{L^{2}(\mu)}\geq\langle\psi(\mathrm{T})\mathrm{S},\mathrm{S}\rangle_{L^{2}(\mu)},$ (22)

### v2

@@ 5 Applications and Experiments
In this section, we first discuss how to verify the relative convexity and smoothness between functionals in practice. Then, we provide some examples of mirror descent and preconditioned gradient descent on different objectives. We refer to Appendix G for more details on the experiments [^1] The code is available at https://github.com/clbonet/Mirror_and_Preconditioned_Gradient_Descent_in_Wasserstein_Space..
@@
To assess relative convexity or smoothness as stated in Definition 3, we need to compare the Bregman divergences along the right curves. When both functionals $\phi$ and $\mathcal{F}$ are of the same type, for example potential (respectively interaction) energies, this property is lifted from the convexity and smoothness on $\mathbb{R}^{d}$ of the underlying potential functions (respectively interaction kernels) to $\mathcal{P}_{2}(\mathbb{R}^{d})$, see Section E.2 for more details. When both $\phi$ and $\mathcal{F}$ are potential energies, the schemes (4) and (10) are equivalent to parallel MD and preconditioned GD since there are no interactions between the particles. The conditions of convergence then coincide with the ones obtained for MD and preconditioned GD on $\mathbb{R}^{d}$ [88, 89]. In other cases, (4) and (10) provide schemes that are novel to the best of our knowledge.
For functionals which are not of the same type, it is less straightforward. Using equivalent notions of convexity (see Proposition 13), we may instead compare their Hessians along the right curves, see Section E.2 for an example between an interaction and a potential energy. We note also that for the particular case of a functional obtained as a sum $\mathcal{F}=\mathcal{G}+\mathcal{H}$ with $\tilde{\mathcal{G}}_{\mu}$ and ~$\tilde{\mathcal{H}}_{\mu}$ convex, since $\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}}=\mathrm{d}_{\tilde{\mathcal{G}}_{\mu}}+\mathrm{d}_{\tilde{\mathcal{H}}_{\mu}}$, $\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}}\geq\max\{\mathrm{d}_{\tilde{\mathcal{G}}_{\mu}},\mathrm{d}_{\tilde{\mathcal{H}}_{\mu}}\}$, and thus $\mathcal{F}$ is 1-convex relative to $\mathcal{G}$ and $\mathcal{H}$. This includes e.g. the KL divergence which is convex relative to the potential and the negative entropy.
MD on interaction energies.
We first focus on minimizing interaction energies $\mathcal{W}(\mu)=\frac{1}{2}\iint W(x-y)\ \mathrm{d}\mu(x)\mathrm{d}\mu(y)$ with kernel $W(z)=\frac{1}{4}\|z\|_{\Sigma^{-1}}^{4}-\frac{1}{2}\|z\|_{\Sigma^{-1}}^{2}$, $\Sigma\in S_{d}^{++}(\mathbb{R})$, whose minimizer is an ellipsoid [27]. Since the Hessian norm of $W$ can be bounded by a polynomial of degree 2, following [88, Section 2], $W$ is $\beta$-smooth relative to $K_{4}(z)=\frac{1}{4}\|z\|_{2}^{4}+\frac{1}{2}\|z\|_{2}^{2}$ with $\beta=4$, and $\mathcal{W}$ is $\beta$-smooth relative to $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\iint K_{4}\big{(}\mathrm{T}(x)-\mathrm{T}(y)\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(y)$. Supposing additionally that the distributions are compactly supported, we can show that $\mathcal{W}$ is smooth relative to the interaction energy with $K_{2}(z)=\frac{1}{2}\|z\|_{2}^{2}$. For ill-conditioned $\Sigma$, i.e. for which the ratio between the largest and smallest eigenvalues is large, the convergence can be slow. Thus, we also propose to use $K_{2}^{\Sigma}(z)=\frac{1}{2}\|z\|_{\Sigma^{-1}}^{2}$ and $K_{4}^{\Sigma}(z)=\frac{1}{4}\|z\|_{\Sigma^{-1}}^{4}+\frac{1}{2}\|z\|_{\Sigma^{-1}}^{2}$. We illustrate these mirror descent schemes on Figure 2 and observe the convergence we expect for the ones taking into account $\Sigma$. In practice, since $\nabla\phi_{\mu}(\mathrm{T})=(\nabla K\star\mathrm{T}_{\#}\mu)\circ\mathrm{T}$, the scheme (5) needs to be approximated using Newton’s algorithm which can be computationally heavy. Using $\phi_{\mu}^{V}(\mathrm{T})=\int V\circ\mathrm{T}\ \mathrm{d}\mu$ with $V=K_{2}^{\Sigma}$, we obtain a more computationally friendly scheme with the same convergence, see Section G.2, but for which the smoothness is trickier to show.
MD on KL.
We now focus on minimizing $\mathcal{F}(\mu)=\int V\mathrm{d}\mu+\mathcal{H}(\mu)$ for $V(x)=\frac{1}{2}x^{T}\Sigma^{-1}x$ with $\Sigma$ possibly ill-conditioned, whose minimizer is the Gaussian $\nu=\mathcal{N}(0,\Sigma)$, and for which Wasserstein gradient descent is slow to converge. We study the MD scheme in (4) with negative entropy $\mathcal{H}$ as the Bregman potential (NEM), and compare it on Figure 2 with the Forward-Backward (FB) scheme studied in [43] and the ideally preconditioned Forward-Backward scheme (PFB) with Bregman potential $\phi_{\mu}^{V}$ (see (116) in Appendix F). For computational purpose, we restrain the minimization in (4) over affine maps, which can be seen as taking the gradient over the submanifold of Gaussians [73, 43]. Starting from $\mathcal{N}(0,\Sigma_{0})$, the distributions stay Gaussian over the flow, and their closed-form is reported in (62) (Section D.3). We note that this might not be the case for the scheme (4), and thus that this scheme does not enter into the framework developed in the previous sections. Nonetheless, it demonstrates the benefits of using different Bregman potentials. We generate 20 Gaussian targets $\nu$ on $\mathbb{R}^{10}$ with $\Sigma=UDU^{T}$, $D$ diagonal and scaled in log space between 1 and 100, and $U$ a uniformly sampled orthogonal matrix, and we report the averaged KL over time. Surprisingly, NEM, which does not require an ideal (and not available in general) preconditioner, is almost as fast to converge as the ideal PFB, and much faster than the FB scheme.
@@
Preconditioned GD for single-cells.
Predicting the response of cells to a perturbation is a central question in biology. In this context, as the measuring process is destructive, feature descriptions of control and treated cells must be dealt with as (unpaired) source $\mu$ and target distributions $\nu$. Following [112], OT theory to recover a mapping $\mathrm{T}$ between these two populations has been used in [21, 23, 22, 39, 48, 122, 69]. Inspired by the recent success of iterative refinement in generative modeling, through diffusion [61, 115] or flow-based models [85, 81], our scheme (1) follows the idea of transporting $\mu$ to $\nu$ via successive and dynamic displacements instead of, directly, with a static map $\bar{\mathrm{T}}$. We model the transition from unperturbed to perturbed states through the (preconditioned) gradient flow of a functional $\mathcal{F}(\mu)=D(\mu,\nu)$ initialized at $\mu_{0}=\mu$, where $D$ is a distributional metric, and predict the perturbed population via $\hat{\mu}=\min_{\mu}\mathcal{F}(\mu)$. We focus on the datasets used in [21], consisting of cell lines analyzed using (i) 4i [58], and (ii) scRNA sequencing [118]. For each profiling technology, the response to respectively (i) 34 and (ii) 9 treatments are provided. As in [21], training is performed in data space for the 4i data and in a latent space learned by the scGen autoencoder [87] for the scRNA data. We use three metrics: the Sliced-Wasserstein distance $\mathrm{SW}_{2}^{2}$ [16], the Sinkhorn divergence $\mathrm{S}_{\varepsilon,2}^{2}$ [50] and the energy distance $\mathrm{ED}$ [105, 60, 59], and we compare the performances when minimizing this functional via preconditioned GD vs. (vanilla) GD. We measure the convergence speed when using a fixed relative tolerance $\textrm{tol}=10^{-3}$, as well as the attained optimal value $\mathcal{F}(\hat{\mu})$. Note that we follow [21] and additionally consider 40% of unseen (test) target cells for evaluation, i.e., for computing $\mathcal{F}(\hat{\mu})=D(\hat{\mu},\nu)$. As preconditioner, we use the one induced by $h^{*}(\*x)=(\|x\|_{2}^{a}+1)^{1/a}-1$ with $a>0$, which is well suited to minimize functionals which grow in $\|x-x^{*}\|^{a/(a-1)}$ near their minimum [119]. We set the step size $\tau=1$ for all the experiments. Then, we tune the parameter $a$ very simply: for a given metric $D$ and a profiling technology, we pick a random treatment and select $a\in\{1.25,1.5,1.75\}$ by grid search, and we generalize the selected $a$ for all the other treatments. Results are described in Figure 3: Preconditioned GD significantly outperforms GD over the 43 datasets, in terms of convergence speed and optimal value $\mathcal{F}(\hat{\mu})$. For instance, for $D=\mathrm{S}_{2,\varepsilon}^{2}$, we converge in 10 times less iterations while providing, on average, a better estimate of the treated population. We also compare our iterative (non parametric) approach with the use of a static (non parametric) map in Section G.4.
@@ 6 Conclusion
Clément Bonet acknowledges the support of the center Hi! PARIS and of ANR PEPR PDE-AI. Adam David gratefully acknowledges funding by the BMBF 01|S20053B project SALE. Pierre-Cyril Aubin-Frankowski was funded by the FWF project P 36344-N. Anna Korba acknowledges the support of ANR-22-CE23-0030.
@@
The appendix is organized as follows. In Appendix A, we discuss related works. In Appendix B and Appendix C, we provide mathematical background respectively on $L^{2}$ and on the Wasserstein space. In Appendix D, we provide complementary results for the mirror descent scheme on the Wasserstein space. In Appendix E, we discuss the relative convexity and smoothness between functionals. In Appendix F, we study the Bregman proximal gradient scheme, deriving a convergence result and closed-form updates for the Gaussian case. In Appendix G, we provide more details on the experiments. In Appendix H, we report the proofs of our results. Finally, auxiliary results are given in Appendix I.
@@ Appendix A Related works
Mirror descent on $\mathbb{R}^{d}$. Mirror descent has been introduced by Nemirovskij and Yudin [92] to solve convex optimization problems. Its convergence has been first studied under the assumption that the objective has a Lipschitz gradient, see e.g. [12]. More recently, Bauschke et al. [11], Lu et al. [88] provided convergence guarantees by assuming relative smoothness and convexity.
Preconditioned gradient descent on $\mathbb{R}^{d}$. The preconditioned gradient descent has first been studied by Maddison et al. [89], providing convergence guarantees under assumptions on the smoothness and convexity of the preconditioner relative to the Legendre transform of the objective. Kim et al. [68] underlined connections with the mirror descent, and introduced an accelerated version of the preconditioned gradient descent. Laude and Patrinos [75], Laude et al. [76] studied a generalized version of this algorithm by minimizing an anisotropic upper bound and supposing anisotropic smoothness of the objective. In particular, their analysis for the descent lemma is also valid for a non-convex smooth objective. Tarmoun et al. [119] also studied preconditioned gradient descent for non-Lipschitz smooth non-convex problems.
@@
The space of probability distributions can be endowed with different metrics. When endowed with the Fisher-Rao metric instead of the Wasserstein distance, the geometry becomes very different. Notably, the shortest path between the two distributions is now a mixture between them. In this situation, the gradient is the first variation. Aubin-Frankowski et al. [9] studied the mirror descent in this space and notably showed connections with the Sinkhorn algorithm when the mirror map and the optimized functionals are KL divergences. Karimi et al. [67] extended the mirror descent algorithm for more general time steps, and notably recovered the “Wasserstein Mirror Flow” proposed by Deb et al. [41] as a special case.
@@ B.1 Differential calculus on $L^{2}(\mu)$
The Fréchet derivative of $\phi$ at $\mathrm{T}\in L^{2}(\mu)$ in the direction $h\in L^{2}(\mu)$, denoted $\delta\phi(\mathrm{T},h)$, is defined implicitly by
@@
If $\phi$ is Fréchet differentiable, then it is also Gâteaux differentiable, and both derivatives agree, i.e. for all $\mathrm{T},h\in L^{2}(\mu)$, $\delta\phi(\mathrm{T},h)=\phi^{\prime}(\mathrm{T})(h)$ [99, Proposition 1.26].
@@ B.2 Convexity on $L^{2}(\mu)$
which is equivalent by [99, Proposition 3.10] with
@@
We now present equivalent definitions of the relative smoothness and relative convexity, which is the equivalent of [88, Proposition 1.1].
@@
    $\psi$ is $\beta$-smooth relative to $\phi$
@@
    $\psi$ is $\alpha$-convex relative to $\phi$
@@
For the rest of the equivalences, we apply [99, Proposition 3.10]. Indeed, $\beta\phi-\psi$ convex is equivalent to
$\forall\mathrm{T},\mathrm{S}\in L^{2}(\mu),\ \langle\nabla(\beta\phi-\psi)(\mathrm{T})-\nabla(\beta\phi-\psi)(\mathrm{S}),\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)}\geq 0\\ \iff\forall\mathrm{T},\mathrm{S}\in L^{2}(\mu),\ \beta\langle\nabla\phi(\mathrm{T})-\nabla\phi(\mathrm{S}),\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)}\geq\langle\nabla\psi(\mathrm{T})-\nabla\psi(\mathrm{S}),\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)},$ (23)
which gives the equivalence between (a2) and (a4). And if $\psi$ and $\phi$ are twice differentiables, it is also equivalent to
$\forall\mathrm{T},\mathrm{S}\in L^{2}(\mu),\ \langle\nabla^{2}(\beta\phi-\psi)(\mathrm{T})\mathrm{S},\mathrm{S}\rangle_{L^{2}(\mu)}\geq 0\\ \iff\forall\mathrm{T},\mathrm{S}\in L^{2}(\mu),\ \beta\langle\nabla^{2}\phi(\mathrm{T})\mathrm{S},\mathrm{S}\rangle_{L^{2}(\mu)}\geq\langle\nabla^{2}\psi(\mathrm{T})\mathrm{S},\mathrm{S}\rangle_{L^{2}(\mu)},$ (24)